OpenStudy (anonymous):

Which related equation can be used to solve the equation? Column A Column B 1. –8.7 = k + 6 2. –6 = k + 8.7 3. 8.7 = k + (–6) 4. 6 = k + 8.7 A. k = –8.7 – 6 B. k = 6 – 8.7 C. k = 8.7 – (–6)

OpenStudy (anonymous):

@dmndlife24 @I_Need_Help_With_Home @Kit_Kat_Nat_<3 @taylor12344

OpenStudy (anonymous):

i am struggleing with this please help

OpenStudy (anonymous):

for each question 1-4 doing them one at a time solve for k make k the subject remember to keep the balance what you do to one side of an equation also do to the other . you have 1 -4 k = whatever now look in column be and find the choice that matches. for example 2. –6 = k + 8.7 k is on the right to get k only -8.7 keep the balance so -8.7 from the left also you get -6 -8.7 = k which matches option A solution 2 A

OpenStudy (anonymous):

@I_Need_Help_With_Home copyed and pasted @triciaal commet